NEW RINGDOWN-DERIVED SPECTROSCOPIC TECHNIQUES FOR THE END OF THE MILLENIUM

\underline{J. J. SCHERER} and A. O’KEEFE, Los Gatos Research 67 E. Evelyn Ave., Suite #3, Mountain View, CA 94041.

Two new, ringdown-based techniques are described which significantly enhance the current capabilities of ringdown spectroscopy. Experimental data is presented for both new methods, with emphasis on routine performance levels for different apparatus configurations.